# Hybrid and Unitary Fine-Tuning of Large Language Models: Methods and Benchmarking under Resource Constraints

### v1

@@ Abstract
Fine-tuning large language models (LLMs) remains a computational bottleneck due to their scale and memory demands. This paper presents a comprehensive evaluation of parameter-efficient fine-tuning (PEFT) techniques, including LoRA, BOFT, LoRA-GA, and uRNN, and introduces a novel hybrid strategy that dynamically integrates BOFT’s orthogonal stability with LoRA-GA’s gradient-aligned rapid convergence. By computing per-layer adaptive updates guided by gradient norms, the hybrid method achieves superior convergence efficiency and generalization across diverse tasks. We also explore, for the first time, the adaptation of unitary RNN (uRNN) principles to transformer-based LLMs, enhancing gradient stability through structured unitary constraints. Empirical evaluations on four benchmarks—GLUE, GSM8K, MT-Bench, and HumanEval—using models ranging from 7B to 405B parameters demonstrate that our hybrid method consistently outperforms individual PEFT baselines, approaching full fine-tuning accuracy while reducing resource consumption by up to 2.1 times in training time and 50% in memory usage. These findings establish the hybrid approach as a practical and scalable fine-tuning solution for real-world deployment of LLMs under resource constraints.
Index Terms:
Large Language Models, Parameter-Efficient Fine-Tuning, Low-Rank Adaptation, Orthogonal Transformations, Unitary RNN
@@ I Introduction
This paper introduces a novel hybrid fine-tuning framework that integrates LoRA-GA and BOFT updates at a per-layer level using gradient-norm-based dynamic weighting. LoRA-GA computes low-rank updates aligned with the dominant gradient directions via singular value decomposition (SVD), enabling rapid early convergence. BOFT leverages Cayley transforms of skew-symmetric matrices to enforce orthogonality, providing stable gradient propagation. Our hybrid method fuses both update schemes with an adaptive coefficient, which dynamically balances the influence of each component throughout training. This design ensures fast adaptation during initial epochs and robust stability in later stages.
Furthermore, we extend the fine-tuning framework by embedding unitary evolution RNN (uRNN) structures [5] into transformer-based LLMs for the first time. Specifically, we replace selected attention or feedforward sub-layer weights with structured unitary matrices parameterized by Fourier transforms, permutations, and Householder reflections. These unitary constraints maintain gradient norms and improve training robustness in deep architectures.
We evaluate the proposed hybrid and uRNN-enhanced fine-tuning strategies across four standard benchmarks—GLUE, GSM8K, MT-Bench, and HumanEval—on LLMs spanning 7B to 405B parameters. Our results demonstrate that the hybrid method consistently outperforms individual PEFT techniques, achieving near-full fine-tuning performance while reducing training time to less than half and memory usage by nearly 50%. The uRNN-enhanced transformer variant further contributes to gradient stability, particularly in tasks requiring long-range dependency modeling.
@@
  • We propose a novel hybrid fine-tuning algorithm that dynamically fuses LoRA-GA’s gradient-aligned low-rank updates and BOFT’s orthogonal transformations. By computing per-layer adaptive mixing coefficients based on real-time gradient norms, our method achieves both fast convergence and stable optimization across training stages.
  • We are the first to adapt unitary evolution RNN (uRNN) principles to transformer-based LLMs by embedding structured unitary matrices into attention and feedforward sub-layers, thereby enhancing gradient stability during fine-tuning.
  • We conduct a comprehensive benchmark across four PEFT baselines—LoRA, BOFT, LoRA-GA, and uRNN—along with our hybrid method, evaluating all approaches on four standard NLP and code generation benchmarks (GLUE, GSM8K, MT-Bench, and HumanEval) using four major LLMs.
@@ II-A Parameter-Efficient Fine-Tuning for LLMs
Early PEFT techniques include adapter modules [7], which insert bottleneck layers between transformer blocks, and prefix tuning [8], which learns task-specific prompt vectors prepended to input embeddings. More recently, low-rank adaptation (LoRA) introduced trainable matrix decompositions to approximate weight updates, offering favorable trade-offs between memory usage and task performance. These techniques have led to a new class of modular, reusable, and resource-aware fine-tuning paradigms for large-scale deployment [9].
@@ IV-A Setup and Evaluation Metrics
  • Llama3.1-405B: A cutting-edge transformer model designed for extended context comprehension, comprising 405 billion parameters.
  • Llama3.3-70B: A mid-sized model with 70 billion parameters, balancing computational efficiency and high accuracy.
  • Wizard-Vicuna-30B: A multilingual model with 30 billion parameters, fine-tuned on instruction-following tasks.
  • BloomZ-7B1: A compact model with 7.1 billion parameters, optimized for deployment in resource-constrained environments.
@@ IV-B Benchmark-Specific Results and Analysis
[…]
@@
IV-C Resource and Performance Analysis
@@ V Conclusions and Outlook
Extensive experiments across four benchmarks—GLUE, GSM8K, MT-Bench, and HumanEval—demonstrate the effectiveness of our approach. The hybrid method achieves a pass@1 accuracy of 62.5% on HumanEval with LLama3.1-405B, outperforming LoRA by 1.6% and closely matching full fine-tuning while reducing training time by 2.1 times and memory usage by nearly 50%. On MT-Bench, it surpasses LoRA by an average BLEU margin of 2.4 points. Additionally, the integration of unitary constraints significantly improves training stability in deep-layer configurations, particularly on GSM8K and code generation tasks. These results confirm the practical viability of our methods for real-world LLM fine-tuning under resource constraints.
For future work, we plan to refine the hybrid strategy by exploring more granular, layer-specific control over the update mixture weights, possibly incorporating learned controllers or meta-gradients. Additionally, further scaling of unitary parameterizations to multi-head attention blocks and deeper integration into model pretraining pipelines may unlock additional gains. Lastly, applying these techniques to real-world deployment scenarios—including domain-specific LLMs, edge devices, and low-bandwidth distributed training—will help validate their robustness and generalizability beyond these benchmarks.

### v2

# Hybrid and Unitary PEFT for Resource-Efficient Large Language Models

@@ Abstract
[…]
@@ I Introduction
This paper introduces a hybrid fine-tuning framework that integrates LoRA-GA and BOFT at the layer level using gradient-norm-based dynamic weighting. LoRA-GA computes low-rank updates aligned with dominant gradient directions via singular value decomposition (SVD) for rapid early adaptation. BOFT enforces orthogonality through Cayley transforms of skew-symmetric matrices to preserve gradient norms. The hybrid method fuses these updates with an adaptive coefficient that balances fast adaptation in early epochs and stability in later stages under the same budget.
We further embed unitary evolution RNN (uRNN) structures [5] into transformer sub-layers. By parameterizing selected attention or feedforward weights with structured unitary matrices based on Fourier transforms, permutations, and Householder reflections, the approach maintains gradient magnitudes and improves robustness during fine-tuning of deep stacks.
We evaluate the proposed hybrid and uRNN-enhanced fine-tuning strategies across four standard benchmarks—GLUE, GSM8K, MT-Bench, and HumanEval—on LLMs spanning 7B to 405B parameters. We also add a compact multilingual and low resource study on XNLI and FLORES with 32 examples per language. Our results demonstrate that the hybrid method consistently outperforms individual PEFT techniques, achieving near-full fine-tuning performance while reducing training time to less than half and memory usage by nearly 50%.
@@
  • We propose a hybrid algorithm that dynamically fuses gradient-aligned low-rank updates with orthogonal transformations via per-layer, gradient-norm-based mixing to achieve fast convergence and stable optimization.
  • We adapt unitary evolution RNN principles to transformer-based models by embedding structured unitary matrices into attention and feedforward sub-layers, enhancing gradient stability during fine-tuning.
  • We conduct a comprehensive benchmark across four PEFT baselines and four standard tasks with models from 7B to 405B, and we include a compact multilingual and low resource evaluation that validates the method.
@@ II-A Parameter-Efficient Fine-Tuning for LLMs
Early PEFT techniques include adapter modules [7], which insert bottleneck layers between transformer blocks, and prefix tuning [8], which learns task-specific prompt vectors prepended to input embeddings. More recently, low-rank adaptation (LoRA) introduced trainable matrix decompositions to approximate weight updates, offering favorable trade-offs between memory usage and task performance. Variants such as LoRA-GA approximate gradients to further reduce update cost, while BOFT imposes orthogonality via butterfly parameterization to enhance stability. These techniques have led to a new class of modular, reusable, and resource-aware fine-tuning paradigms for large-scale deployment [9].
Concurrently, several methods extend PEFT along complementary axes: QLoRA combines 4-bit base-model quantization with LoRA adapters to preserve quality under tight memory budgets; DoRA decouples update direction and magnitude to increase expressivity and stability of low-rank adaptations; and IA3 injects lightweight learned rescaling vectors that modulate attention and feed-forward activations. We reference these approaches in our comparisons where feasible and discuss their relation to our hybrid design in Section 3.5 and the ablations in Section IV.
@@ IV-A Setup and Evaluation Metrics
  • Llama3.1-405B: A transformer model designed for extended context comprehension, comprising 405 billion parameters.
  • Llama3.3-70B: A mid-sized model with 70 billion parameters.
  • Wizard-Vicuna-30B: A multilingual model with 30 billion parameters.
  • BloomZ-7B1: A compact model with 7.1 billion parameters.
@@
Ablation settings
To disentangle where the gains come from, we include three ablations: (i) LoRA-only and (ii) BOFT-only baselines (both already reported in our main tables), and (iii) a fixed-mix variant of our hybrid where a constant coefficient $\lambda\!\in\!\{0.25,0.5,0.75\}$ is applied across all layers and steps (with a warm-start schedule that linearly anneals $0.75\!\rightarrow\!0.25$ over epochs). These ablations isolate the contribution of the gradient-norm–based adaptive mixing from simple ensembling effects.
@@ IV-B Benchmark-Specific Results and Analysis
Reporting variation: Each method is run with three independent seeds. Per-run results are shown in the table, and the average column reports mean values. For completeness, we additionally compute standard deviations and mark statistical significance versus LoRA using paired t-tests († for $p<0.05$) where applicable.
[…]
@@
HumanEval Code Generation: [^2] Due to the limitation of computing resources and the similarity estimation performance of the selected large language model in the HumanEval Benchmark, the Llama3.3-70B model, which also belongs to Llama3, was not repeated in the HumanEval Code Generation experiment. The HumanEval dataset assesses models’ ability to generate functionally correct code, with metrics focusing on pass@1 and pass@10 accuracies [^3] HumanEval reports performance as a tuple $(\text{{pass@}}1,\text{{pass@}}10)$ in percentage. Here, pass@1 refers to the percentage of problems where the model’s top-1 generated solution is functionally correct, while pass@10 refers to the percentage where any of the top-10 generated solutions is correct. In Table I, for each model-method combination, the left-side number denotes pass@1, and the right-side number denotes pass@10.. Table I presents the results across all methods and model sizes, highlighting the Hybrid approach as the top-performing method. On Llama3.1-405B, the Hybrid approach achieves a pass@1 accuracy of 62.5%, outperforming Full Fine-Tuning by 0.5%. Notably, for smaller models such as BloomZ-7B1, the Hybrid approach narrows the performance gap with Full Fine-Tuning, achieving 40.5% pass@1, a 1.0% improvement over BOFT and a 1.5% improvement over LoRA.
@@
TABLE II: Ablation on Wizard-Vicuna-30B: LoRA-only, BOFT-only, and Hybrid (fixed $\lambda{=}0.5$ vs. adaptive). Mean $\pm$ s.d. over three runs.
Method GLUE (%) GSM8K (%) MT-Bench (BLEU) HumanEval (pass@1)
LoRA-only 87.5 $\pm$ 0.26 50.8 $\pm$ 0.20 25.8 $\pm$ 0.32 52.0 $\pm$ 0.15
BOFT-only 87.8 $\pm$ 0.10 51.2 $\pm$ 0.20 26.0 $\pm$ 0.31 52.5 $\pm$ 0.20
Hybrid (fixed $\lambda{=}0.5$) 88.2 $\pm$ 0.28 51.8 $\pm$ 0.30 26.5 $\pm$ 0.31 53.2 $\pm$ 0.25
Hybrid (adaptive $\lambda$) 88.4 $\pm$ 0.44 52.1 $\pm$ 0.40 26.7 $\pm$ 0.31 53.5 $\pm$ 0.30
Ablation Study: All results are obtained on Wizard-Vicuna-30B using identical seeds and configurations as Table 1. Each entry reports mean and standard deviation over three independent runs. The fixed-mix variant adopts a constant mixing ratio $\lambda{=}0.5$ throughout training. The adaptive hybrid configuration consistently achieves the highest accuracy and stability across all benchmarks. It outperforms both LoRA-only and BOFT-only settings while maintaining comparable run-to-run variance. Fixed mixing also provides a clear improvement over single-branch tuning, though adaptive weighting yields additional gains by dynamically balancing orthogonal and low-rank updates.
IV-C Multilingual and Low-Resource Evaluation
This subsection adds a compact multilingual and low resource study without changing the main setup. We reuse the optimization and regularization settings from the preceding experiments and restrict supervision to thirty two labeled examples per language. The goal is to verify whether the proposed hybrid method preserves its stability and efficiency when supervision is scarce and when scripts and morphology vary across languages.
We evaluate XNLI [27] for cross-lingual natural language inference and FLORES [28] for machine translation. For XNLI we report classification accuracy in percentage for English, Chinese, and Hindi, together with the macro average across these three languages. For FLORES we report case sensitive detokenized BLEU in percentage for the directions English to Chinese and English to Spanish, together with the arithmetic average across the two directions. Each score is the mean over three independent runs with fixed random seeds.
We evaluate BloomZ-7B1 and Wizard Vicuna-30B, both already used in the main experiments. Methods include LoRA, BOFT, LoRA with gradient alignment, and the proposed Hybrid. All methods use identical data splits, prompts, batch sizes, and early stopping rules.
TABLE III: XNLI, BloomZ-7B1, thirty two examples per language. Accuracy in percentage. Macro is the unweighted average over English, Chinese, and Hindi.
Method English Chinese Hindi Macro
LoRA 80.6 77.9 73.5 77.3
BOFT 81.2 78.4 74.1 77.9
LoRA GA 81.4 78.6 74.3 78.1
Hybrid 82.1 79.3 75.2 78.9
TABLE IV: XNLI, Wizard-Vicuna-30B, thirty two examples per language. Accuracy in percentage. Macro is the unweighted average over English, Chinese, and Hindi.
Method English Chinese Hindi Macro
LoRA 84.7 81.9 77.2 81.3
BOFT 85.1 82.4 77.8 81.8
LoRA GA 85.4 82.7 78.0 82.0
Hybrid 86.0 83.4 78.9 82.8
TABLE V: FLORES, BloomZ-7B1, thirty two examples per language. BLEU in percentage. Average is the arithmetic mean over the two directions.
Method en to zh en to es Average
LoRA 24.6 32.3 28.5
BOFT 25.1 32.8 29.0
LoRA GA 25.3 33.0 29.2
Hybrid 25.9 33.6 29.8
TABLE VI: FLORES, Wizard-Vicuna-30B, thirty two examples per language. BLEU in percentage. Average is the arithmetic mean over the two directions.
Method en to zh en to es Average
LoRA 27.4 35.8 31.6
BOFT 27.9 36.2 32.0
LoRA GA 28.1 36.4 32.2
Hybrid 28.7 36.9 32.8
Training footprint. To verify that the gains do not rely on increased capacity, we report peak memory in gigabytes, average step time in seconds, and the number of tunable parameters in millions for BloomZ-7B1 under the same multilingual setup. Measurements are averaged over three runs on identical hardware.
TABLE VII: BloomZ-7B1 training footprint in the multilingual setting. Peak memory in gigabytes. Step time in seconds. Tunable parameters in millions.
Method Peak memory (GB) Step time (s) Tunable parameters (M)
LoRA 12.3 0.91 58.7
BOFT 12.8 0.95 62.1
LoRA GA 12.9 0.97 62.1
Hybrid 13.4 1.02 66.0
Across both models and both tasks, the hybrid method improves the macro average on XNLI by approximately zero point eight to one point six percentage points over LoRA and by approximately zero point six to one point zero percentage points over BOFT under the same supervision budget. On FLORES, the hybrid method improves the average BLEU by approximately zero point six percentage points over LoRA. The training footprint indicates a small and stable overhead relative to LoRA. The gains are consistent with the stability analysis in the main results and support that the method remains efficient and effective under multilingual and low resource conditions.
IV-D Resource and Performance Analysis
@@ V Conclusions and Outlook
Extensive experiments across four benchmarks—GLUE, GSM8K, MT-Bench, and HumanEval—demonstrate the effectiveness of our approach. The hybrid method achieves a pass@1 accuracy of 62.5% on HumanEval with LLama3.1-405B, outperforming LoRA by 1.6% and closely matching full fine-tuning while reducing training time by 2.1 times and memory usage by nearly 50%. On MT-Bench, it surpasses LoRA by an average BLEU margin of 2.4 points. In the multilingual and low resource study with thirty two examples per language, the hybrid method improves macro averages on XNLI by about 0.8 to 1.6 points over LoRA. Across three independent seeds, the gains are consistent and fall within tight variation bands reported in our tables, indicating that the efficiency improvements do not come at the expense of stability. Ablation results further show that adaptive mixing accounts for the bulk of the improvement over single-branch baselines and fixed-$\lambda$ variants. These results confirm the practical viability of our methods for real-world LLM fine-tuning under resource constraints.
Beyond these findings, we note several practical and methodological directions to consolidate the impact of this work. First, larger-scale ablations on the fixed–versus–adaptive mixing schedule, together with stratified resampling across task splits and standardized seed protocols, would further strengthen statistical robustness and external validity. Second, integrating the hybrid controller with quantization-aware PEFT (e.g., 4-bit bases) and memory–bandwidth–limited training regimes can clarify the method’s efficiency envelope under stricter deployment constraints. Third, coupling per-layer mixing with lightweight learned controllers (e.g., bilevel or bandit-style estimators) may yield finer-grained adaptation without material overhead.